### v3

# Efficient Context Selection for Long-Context QA: No Tuning, No Iteration, Just Adaptive‑$k$

Abstract
Retrieval-augmented generation (RAG) and long-context language models (LCLMs) both address context limitations of LLMs in open-domain question answering (QA). However, optimal external context to retrieve remains an open problem: fixing the retrieval size risks either wasting tokens or omitting key evidence. Existing adaptive methods like Self-RAG and Self-Route rely on iterative LLM prompting and perform well on factoid QA, but struggle with aggregation QA, where the optimal context size is both unknown and variable. We present Adaptive‑$k$ retrieval, a simple and effective single-pass method that adaptively selects the number of passages based on the distribution of the similarity scores between the query and the candidate passages. It does not require model fine-tuning, extra LLM inferences or changes to existing retriever–reader pipelines. On both factoid and aggregation QA benchmarks, Adaptive‑$k$ matches or outperforms fixed‑$k$ baselines while using up to 10× fewer tokens than full-context input, yet still retrieves 70% of relevant passages. It improves accuracy across five LCLMs and two embedding models, highlighting that dynamically adjusting context size leads to more efficient and accurate QA. [^1] The code is available at https://github.com/megagonlabs/adaptive-k-retrieval.
Efficient Context Selection for Long-Context QA: No Tuning, No Iteration, Just Adaptive‑$k$
Chihiro Taguchi thanks: Work done during internship at Megagon Labs. University of Notre Dame ctaguchi@nd.edu          Seiji Maekawa Megagon Labs seiji@megagon.ai          Nikita Bhutani Megagon labs nikita@megagon.ai
1 Introduction
Despite remarkable progress in LLMs, efficiently incorporating external knowledge during inference for long or dynamic contexts remains a key challenge. Two major paradigms have emerged to address this: long-context language models (LCLMs), which extend the model’s context window to directly ingest more information, and retrieval-augmented generation (RAG), which retrieves relevant documents from an external corpus to condition the generation. While these approaches are sometimes presented as alternatives Li et al. (2024a); Yu et al. (2024), recent studies highlight their complementary nature Li et al. (2024b).
A central bottleneck in both paradigms is determining how much context to include. Fixed-size retrieval budgets (e.g., top-$k$ retrieval) are suboptimal, because they either retrieve too little and risk omitting key evidence, or retrieve too much, which can overwhelm the model, increase latency and costs, and degrade performance Yu et al. (2024); Leng et al. (2024); Jin et al. (2024). As Yang (2024) observes, the challenge in long-context reasoning lies not only in document length but also in how relevant information is distributed and duplicated within the context. Crucially, query type plays a major role: factoid questions may need only a few targeted facts, while aggregation queries Maekawa et al. (2025) often require reasoning based on information from multiple evidence spans. This variability makes fixed-$k$ retrieval suboptimal for complex tasks.
To address this, several hybrid and adaptive retrieval methods such as Self-RAG Asai et al. (2023), Adaptive-RAG Jeong et al. (2024), and Dynamic context cutoff Xie et al. (2025) have been proposed, which estimate retrieval depth via iterative prompting, each time fetching a fixed number of documents. However, they assume white-box access to the LLM: Self-RAG requires fine-tuning the LLM, while dynamic context cutoff depends on access to internal KV cache states. This makes them incompatible with closed-source or API-based LLMs. While effective on factoid-style questions, they also face significant limitations in terms of scalability, latency, and deployment flexibility. Although Self-Route Li et al. (2024b) offers a more modular solution, it still relies on a fixed retrieval size and lacks the ability to adapt to varying information needs across queries and context documents. This motivates our core research question: How can we estimate the optimal number of passages to retrieve for a given query and set of context documents, without supervision or iterative prompting?
Plug-and-Play via API Retrieval Amount Variability Single Retrieval Operation
No RAG (LCLM) No Retrieval
RAG (traditional)
Self-RAG Asai et al. (2023)
Adaptive-RAG Jeong et al. (2024)
Self-Route Li et al. (2024b)
LC-Boost Qian et al. (2024)
Dynamic context cutoff Xie et al. (2025)
Adaptive-$k$ RAG (ours)
Table 1: The comparison of previously proposed approaches as enhanced RAG. Plug-and-Play via API refers to whether the approach can be easily plugged in to various LLM pipelines. Retrieval Amount Variability refers to whether the system can flexibly change the retrieval amount depending on different queries and context. Single Retrieval Operation refers to whether the retrieval is performed in a single step or in multiple steps.
To address this question, we introduce Adaptive-$k$ retrieval, a simple yet effective plug‑and‑play method for dynamically selecting a query- and context-specific number of documents in a single retrieval pass. Our approach relies on analyzing the distribution of similarity scores between a query and candidate documents. By identifying the largest gap in the sorted similarity distribution, it estimates an optimal cutoff point, retrieving the top-$k$ documents before the gap. Unlike prior adaptive retrieval methods, Adaptive‑$k$ requires no model fine-tuning, no access to internal components and no iterative prompting. It is fully modular, allowing seamless integration with existing retriever–reader pipeline and compatibility with black-box LLMs. By relying solely on the distributional structure of similarity scores, Adaptive‑$k$ adjusts the retrieval size on a per-query basis. This simple yet principled strategy leads to significant reductions in input length and inference cost, while maintaining or even improving the answer quality across both factoid and aggregation-style QA tasks. We compare Adaptive-$k$ retrieval to prior approaches in Table 1.
We evaluate Adaptive-$k$ on both factoid and aggregation-style QA tasks across multiple LCLMs and embedding models. Our experiments span two representative long-context benchmarks: HELMET Yen et al. (2025), which includes factoid QA tasks with up to 128k-token contexts, and HoloBench Maekawa et al. (2025), which focuses on aggregation-style queries. Our results show that on aggregation-QA, Adaptive-$k$ outperforms Self-Route by up to +9 points in answer accuracy on high-information tasks. It consistently maintains $\sim$70% context recall and reduces token usage by 2× to 10× compared to full-context baselines. On factoid QA, Adaptive-$k$ matches or exceeds the accuracy of fixed-size retrieval with up to 99% reduction in input tokens, effectively pruning irrelevant content. These findings highlight the importance of query-specific context sizing and establish Adaptive-$k$ as a simple, robust, and efficient alternative to more complex adaptive retrieval strategies.
In summary, our key contributions are:
  • We propose Adaptive-$k$, a simple yet effective plug-and-play method for adaptive document retrieval that dynamically adjusts context size based on similarity distribution statistics.
  • Adaptive-$k$ achieves higher accuracy than prior methods and up to 99% token reduction on factoid and aggregation QA against LCLMs with full context.
  • We show that no single fixed-size retrieval strategy fits all settings. In contrast, Adaptive-$k$ shows robust performance across multiple LLMs, embedding models and benchmarks.
2 Related Work
RAG and LCLMs are two prominent paradigms for equipping LLMs with external knowledge. Recent studies show that LCLMs can match or outperform RAG in certain QA tasks Li et al. (2024a); Yu et al. (2024), yet the two methods are fundamentally complementary.
Several approaches have been proposed to leverage both the strengths of RAG and LCLMs with flexible retrieval strategies. Self-RAG Asai et al. (2023) trains an LLM to generate reflection tokens that enable retrieval on the fly, so that the LLM can determine whether it needs any additional document by itself. Self-Route Li et al. (2024b) asks an LLM whether it can answer the query with the retrieved context; if not, the LLM is given the full context. Adaptive-RAG Jeong et al. (2024) uses a workflow that iteratively asks an LLM whether it can answer the given query with the retrieved context. LC-Boost Qian et al. (2024) enables short-context LLMs to tackle long-context tasks by first identifying relevant information, then reasoning over it, without needing extended context windows or fine-tuning.
While effective in controlled settings, these methods often rely on white-box access to the LLM, fine-tuning, or multiple LLM inferences. Existing research has highlighted key limitations in RAG systems, particularly in terms of cost, modularity, and retrieval granularity. However, prior methods typically address these issues in isolation, and to our knowledge, no single approach has tackled all three challenges holistically. Our method is the first to offer a unified solution that is cost-efficient, modular, and capable of adaptive, query-specific retrieval in a single pass.
Cost.
High-quality inference often comes with high token usage, energy consumption, and latency Li et al. (2024b); Qian et al. (2024), underscoring the need for more cost-effective alternatives.
Modularity.
Modularity is crucial for real-world deployment Wang et al. (2024), but many existing methods require fine-tuning or training the LLM itself. This tight coupling reduces compatibility with API-based or closed-source models, limiting practical applicability.
Retrieval granularity.
Aggregation-type queries often require comprehensive evidence and holistic understanding. For example, answering “Which colleges in California have over 10,000 students?” demands access to the full set of relevant entries. Fixed-size or iterative retrieval methods struggle with such cases, as they cannot dynamically adjust retrieval depth based on query complexity.
3 Method
This section details our approach to adaptive retrieval, grounded in the analysis of similarity score patterns to determine retrieval sizes adaptively based on the query and the context. We first review the standard RAG retrieval process, then present our methodology to identify the optimal threshold in similarity distributions to efficiently select relevant documents.
3.1 Retrieval in vanilla RAG
RAG consists of two steps: retrieval and generation. Given a query $q$ and $N$ context documents $C=\{c_{i}\}_{1}^{N}$, the retriever module identifies top-$k$ semantically similar context documents $C^{\prime}$. Modern RAG approaches convert the query and the context documents in natural language into the query embedding $\bm{q}\in\mathbb{R}^{d}$ and context embeddings $\bm{C}\in\mathbb{R}^{N\times d}$. Similarity scores $\bm{s}\in\mathbb{R}^{N}$ are then computed to quantify relevance, commonly using cosine similarity:
[…]
RAG typically retrieves a fixed number of top-$k$ documents (or tokens) based on the practitioner’s choice. This fixed retrieval size is simple and modular but may result in inefficient token usage, either retrieving irrelevant documents or missing critical information, especially when the amount of relevant context varies depending on the provided context documents and the query type.
Figure 1: Example distributions of sorted cosine similarities from the long-context version of HotpotQA Yang et al. (2018) included in HELMET Yen et al. (2025) with 1,000 context documents (top) and HoloBench Maekawa et al. (2025) with 10% relevant information amount (bottom). BAAI’s bge-large-en-v1.5 is used as the embedding model.
[…]
Figure 2: The proposed method in the RAG workflow. The method chooses the threshold k for retrieval based on a large gap in the sorted similarity score distribution.
3.2 Toward efficient adaptive retrieval
Design motivation and principles.
While vanilla RAG offers modularity and straightforward integration, its fixed retrieval size limits performance and efficiency in scenarios where the quantity of relevant context varies unpredictably such as in aggregation QA in the HoloBench benchmark Maekawa et al. (2025). To address these limitations, we aim to design an adaptive retrieval mechanism that: (1) operates independently of the underlying inference model and requires no additional training or fine-tuning (Plug-and-Play), (2) flexibly controls the retrieval amount for each query, avoiding both wasting tokens and omitting key evidence (Retrieval Amount Variability), and (3) operates in a single pass without requiring iterative LLM calls (Single Retrieval Operation).
Preliminary analysis.
To ground our design in empirical evidence, we conduct an in-depth analysis of the distributional patterns of cosine similarity scores between queries and candidate documents, which, crucially, are inference model-agnostic signals. This preliminary analysis reveals distinct distributional characteristics that inform our adaptive retrieval strategy.
As shown in Figure 1, for factoid QA tasks such as HotpotQA, the sorted similarity scores typically exhibit a pronounced gap separating a cluster of highly relevant documents from the rest, suggesting a natural threshold for retrieval. In contrast, aggregation tasks (e.g., HoloBench) show more irregular patterns, with gaps dispersed throughout the distribution – reflecting the variable spread of relevant information. In the bottom example in Figure 1, the 100k-token context is generated such that 10% of it is information relevant to the query. Indeed, the large gaps are observed around the top 5% to 20% context, aligning with our expectations.
These insights lead to the hypothesis that the largest gap in sorted similarity scores corresponds to the boundary between relevant and irrelevant documents, thus providing a data-driven criterion for adaptive retrieval size selection.
3.3 Proposed method
Building on these observations, we formalize an algorithm that adaptively estimates the retrieval threshold $k$ by identifying the position of the steepest drop in the similarity score distribution. The method proceeds as follows: Compute the cosine similarities $\bm{s}$ of the query $\bm{q}$ and context documents $\bm{C}$. Sort the scores in descending order. Compute their first discrete differences $\bm{g}$ and choose the index $k$ where the similarity drop is the largest. Figure 2 depicts this process within the RAG workflow. Under the assumption that the embeddings of documents are precomputed, the time complexity of this algorithm is $\mathcal{O}(n\log n)$. The algorithm is described in Algorithm 1.
Algorithm 1 Adaptive $k$ Estimation via Largest Similarity Gap
$q$, $C$, Embedder($\cdot$), Similarity($\cdot$)
Estimated $k$ such that the largest similarity drop occurs before the $k$-th item
$\bm{q}\leftarrow\texttt{Embedder}(q)$
$\bm{C}\leftarrow\texttt{Embedder}(C)$ $\triangleright$ Precomputed
$\bm{s}\leftarrow\texttt{Similarity}(\bm{q},\bm{C})$
Sort $\bm{s}$ in descending order
$\bm{g}\leftarrow\texttt{array}()$ $\triangleright$ For storing the gap
for $i=0$ to $|\bm{s}|-2$ do
  Append $\bm{s}[i]-\bm{s}[i+1]$ to $\bm{g}$
end for
$k\leftarrow\arg\max(\bm{g})$ $\triangleright$ Index at the largest gap
return $k$
In practice, while determining the threshold $k$ based on the largest similarity gap is effective, a naïve implementation might miss relevant documents located immediately beyond the identified threshold. To address this, we incorporate a small fixed buffer, retrieving an additional $B$ documents after the $k$-th document. In our experiments, we set $B=5$. Furthermore, as depicted in Figure 1, the largest gap may occasionally manifest among the least relevant documents, leading to the retrieval of an excessively large portion of the context. To avoid this and align with our focus on retrieval from extremely long contexts, we restrict the search for the largest gap to the top 90% of documents sorted by their similarity scores.
4 Experimental setup
In our experiments, we aim to answer the following research questions:
  • How does the proposed adaptive-$k$ method compare to other modular retrieval approaches on aggregation tasks with varying amounts of relevant context?
  • How does performance of Adaptive-$k$ vary across factoid QA and aggregation QA tasks?
  • How does the performance gain from Adaptive‑$k$ retrieval vary across LLMs?
  • How do different embedding models influence the performance of Adaptive-$k$?
To answer these questions, we employ the experimental settings detailed below.
4.1 Dataset
For testing on factoid QA tasks, we use HotpotQA Yang et al. (2018), Natural Questions (NQ) Kwiatkowski et al. (2019), and TriviaQA Joshi et al. (2017), as curated by HELMET Yen et al. (2025) for long-context benchmarking with 128k input tokens. Due to the high computational cost of long-context inference, we evaluate on a subset of 100 examples per dataset.
For aggregation tasks, we employ HoloBench Maekawa et al. (2025), which provides 90 evaluation samples. HoloBench allows control over both total context size and the amount of information relevant to the query. We fix the total context to 100k tokens and evaluate under varying levels of relevant information, with $\texttt{info\_amount}=$ {5000, 10000, 25000, 50000} tokens.
4.2 Models
Retriever.
We test our method on small, medium, and large embedding models: Meta’s contriver-msmarco [^2] https://huggingface.co/facebook/contriever-msmarco (Izacard et al., 2021) with 109M params, BAAI’s bge-en-large-v1.5333https://huggingface.co/BAAI/bge-large-en-v1.5 Xiao et al. (2023) with 335M params, and Alibaba NLP’s gte-Qwen2-1.5B-instruct [^4] https://huggingface.co/Alibaba-NLP/gte-Qwen2-1.5B-instruct Li et al. (2023) with 1.78B params.
Reader.
We use five closed and open models: GPT-4o-mini, GPT-4o OpenAI et al. (2024), Gemini-2.5-Flash Team et al. (2024), Llama4-Scout, and Llama4-Maverick Touvron et al. (2023). The model details are provided in Appendix A.2.
4.3 Compared methods
We compare the proposed adaptive-$k$ method against zero-shot LLMs (without context), LLMs with full context, and Self-Route Li et al. (2024b), which is another modular retrieval method with a single retrieval step. In Self-Route, fixed top 5k tokens are retrieved for the first inference step. We also show the results of the fixed-$n$ retrieval method with varying numbers of tokens $n$ as performance references. Specifically, we run experiments with $n\in$ {1000, 5000, 10000, 25000, 50000} and regard the best-performing setting as the oracle. In this way, we can compare the performance of adaptive-$k$ against the best possible score of the fixed retrieval method.
(a) $\texttt{info\_amount}=10000$.
fixed-25k (oracle)full-contextSelf-Routeadaptive-k (ours)020204040020k40k60k80k100kScore# Input Tokens
(b) $\texttt{info\_amount}=25000$.
fixed-25k (oracle)full-contextSelf-Routeadaptive-k (ours)020204040020k40k60k80k100kScore# Input Tokens
(c) $\texttt{info\_amount}=50000$.
fixed-50k (oracle)full-contextSelf-Routeadaptive-k (ours)020204040020k40k60k80k100kScore# Input Tokens
Figure 3: The results with different amounts of relevant information in the HoloBench tasks. The best-performing fixed-n setting is chosen as the oracle.   is for performance improvement, and   for the number of input tokens.
4.4 Metrics
To evaluate the retrieval performance, context recall Ru et al. (2024) is computed, which represents how much of the relevant context documents were able to be retrieved. For the evaluation of generation performance, we use substring exact match (SubEM) for HotpotQA, NQ, and TriviaQA, and LLM-as-a-judge for HoloBench, following the metrics used in their original implementation in HELMET Yen et al. (2025) and HoloBench, respectively. LLM-as-a-judge evaluates whether the generated answer contains a correct mention of the gold answer, assigning a score of 1 if it finds a correct mention, 0.5 for a partially correct mention, and 0 otherwise. For the judge model, GPT-4o-mini is used. To evaluate the inference cost, we count the number of input and output tokens, assuming that the financial cost on the user’s end and energy consumption depends on the amount of tokens Husom et al. (2024).
5 Results
This section provides the results of the experiments with a focus on different task types, reader models, and embedding models. For the full results, see Appendix A.3.
5.1 Aggregation-type QA
Figure 3 shows GPT-4o’s results in the HoloBench tasks where each task is designed to contain different amounts of relevant information (info_amount: 10k, 25k, 50k tokens) in the context. It can be observed that our Adaptive-$k$ method constantly outperforms Self-Route. The performance improvements of Adaptive-$k$ are particularly notable when the amount of relevant information in the context is high. Also, our method flexibly increases the amount of retrieved context chunks when there is a higher amount of relevant information in the entire context. In contrast, Self-Route tends to underestimate the amount of relevant context and jump to a conclusion that the LLM can answer the query with the 5k-token context retrieved in the first round, leading to lower performance in a high amount of relevant information.
This contrast is also reflected in the context recall scores. As shown in Table 2, Adaptive-$k$ consistently achieves a context recall score of approximately 70 across varying levels of relevant information, indicating that it retrieves approximately 70% of truly relevant chunks regardless of their proportion in the full context. The contrast is even more pronounced when compared to context recall of Self-Route, with Adaptive-$k$ achieving more than three times higher context recall.
info5k info10k info25k info50k
Self-Route 65.79 45.04 30.42 21.54
Adaptive-$k$ 75.74 68.54 66.16 67.43
fixed-1k 12.05 6.53 2.77 1.47
fixed-5k 51.92 31.77 14.06 7.54
fixed-10k 66.68 59.10 28.80 15.39
fixed-25k 78.48 78.18 68.13 39.55
fixed-50k 86.79 87.34 86.88 76.90
Table 2: A comparison of the context recall scores across different relevant information amounts in the HoloBench tasks. The query and contexts are embedded by bge-large-en-v1.5. The scores compared are Self-Route and Adaptive-$k$, as well as the results of fixed-$n$ token retrieval as references.
5.2 Factoid-type QA
Figure 4 shows the comparison of Adaptive-$k$ against the zero-shot setting, fixed 1k-token retrieval, full context, and Self-Route. All methods are implemented using GPT-4o. Our method achieves a 99% reduction in input cost compared to the full context input, and a 90% reduction compared to Self-Route. Since users generally lack prior knowledge of the optimal retrieval size, Adaptive-$k$ successfully reduces the cost while improving the generation quality compared to zero-shot question answering.
[…]
Figure 4: A performance comparison of our proposed method (Adaptive-$k$) in the factoid QA tasks against existing methods. The embedding model is bge-large-en-v1.5, and the reader model is GPT-4o.   is for the SubEM scores, and   for the number of input tokens.
5.3 Comparison across LLMs
(a) GPT-4o-mini.
fixed-25k (oracle)full-contextSelf-Routeadaptive-k (ours)0202040406060020k40k60k80k100kScore# Input Tokens
(b) GPT-4o.
fixed-25k (oracle)full-contextSelf-Routeadaptive-k (ours)0202040406060020k40k60k80k100kScore# Input Tokens
(c) Gemini-2.5-Flash.
fixed-50k (oracle)full-contextSelf-Routeadaptive-k (ours)0202040406060020k40k60k80k100kScore# Input Tokens
(d) Llama4-Scout.
fixed-5k (oracle)full-contextSelf-Routeadaptive-k (ours)0202040406060020k40k60k80k100kScore# Input Tokens
(e) Llama4-Maverick.
fixed-50k (oracle)full-contextSelf-Routeadaptive-k (ours)0202040406060020k40k60k80k100kScore# Input Tokens
Figure 5: A performance comparison across the different reader models in the HoloBench task. The emnbedding model is bge-large-en-v1.5.   is for performance improvement, and   for the number of input tokens.
Since our methods only modify the retriever module, the retrieved documents to be fed into an LLM’s prompt remain the same across different LLMs. However, we observe that its effectiveness varies notably by model. Figure 5 shows the average score improvements and input token counts across different relevant information settings in HoloBench. Larger high-performance LLMs such as GPT-4o (Figure 5(b)), Gemini-2.5-Flash (Figure 5(c)), and Llama4-Maverick (Figure 5(e)) show substantial gains from Adaptive-$k$ retrieval compared to Self-Route. In contrast, smaller LLMs such as GPT-4o-mini (Figure 5(a)) and Llama4-Scout (Figure 5(d)) exhibit more modest improvements. Nonetheless, even for smaller models, Adaptive-$k$ effectively reduces context length while maintaining performance close to the full-context and oracle fixed-$n$ baselines.
5.4 Embedding bottleneck
Method Contriever BGE GTE
HotpotQA Self-Route 77.50 90.83 25.83
adaptive-k 49.17 70.83 5.50
NQ Self-Route 51.19 51.90 20.67
adaptive-k 26.28 27.20 2.85
TriviaQA Self-Route 41.49 46.52 10.20
adaptive-k 25.02 31.21 3.00
HoloBench-5k Self-Route 60.57 65.79 65.18
adaptive-k 72.11 75.74 82.20
HoloBench-10k Self-Route 41.83 45.04 45.87
adaptive-k 70.85 68.54 78.99
HoloBench-25k Self-Route 29.67 30.42 31.02
adaptive-k 68.67 66.16 76.47
HoloBench-50k Self-Route 21.27 21.54 21.90
adaptive-k 67.16 67.43 72.54
Table 3: A comparison of the context recall scores across tasks between Contriever (contriever-msmarco), BGE (bge-large-en-v1.5), and GTE (gte-Qwen2-1.5B-instruct).
We observed that the effectiveness of our adaptive method is sensitive to the choice of embedding model. As shown in Table 3, the embeddings by bge-large-en-v1.5, gte-Qwen2-1.5B-instruct, and contriever-msmarco have different strengths depending on the task. In factoid QA tasks, BGE embeddings consistently yield higher context recall than GTE, whereas GTE performs better on HoloBench. The underlying cause remains unclear, but we identify a few potential factors: (1) Context chunk length: the factoid QA tasks in HELMET generally have a longer context chunk length (up to $\sim$100 tokens) than HoloBench ($\sim$40 tokens); (2) Chunking scheme Zhong et al. (2025): while the context chunks in HoloBench contain well-formed natural-language sentences, those in the factoid QA tasks often contain mid-sentence breaks; (3) Training scheme: differences in pretraining corpora and formatting may lead to divergent performance across embedding models. Overall, choosing the right embedding model is critical for ensuring RAG effectiveness. For general use, we recommend bge-large-en-v1.5 for Adaptive-$k$ due to its strong and consistent performance across settings.
5.5 Limitation of fixed retrieval
While fixed-$n$ retrieval occasionally outperforms Adaptive-$k$ method, it requires prior knowledge of the optimal $n$, which is difficult to estimate in practice. Our results show that the best-performing $n$ varies across task types, query types, embedding models, reader models, and the distribution of relevant information. In contrast, Adaptive-$k$ is able to dynamically adjust the retrieval amount based on the query and context chunks, eliminating the need for manual tuning. This not only removes the burden and risk of heuristically selecting an $n$ but also provides a more robust and generalizable solution across a wide range of scenarios, especially in cases where the relevant context size is highly variable or unknown a priori.
In addition, Tables 4 and 5 demonstrate that the Adaptive-$k$ method effectively estimates the threshold of relevant contexts across different relevant context sizes. These tables report the average absolute difference between the estimated threshold ($k$) and the true $k$-value, referred to as diff-$k$ henceforth, across various datasets and embeddings. The true $k$-value is defined as the index of the last relevant context chunk in the list of context chunk embeddings sorted by similarity; in other words, diff-$k$ is the smallest value of $k$ that achieves 100% recall. Table 4 presents the diff-$k$ values for the HoloBench task with 50k tokens of relevant context. The diff-$k$ values for Adaptive-$k$ are substantially lower than those of the Self-Route baseline and closely match those of the fixed 50k-token retrieval, which is treated as an oracle. Similarly, the diff-$k$ values for Adaptive-$k$ in the HotpotQA task (Table 5) consistently approximate the oracle retrieval results, indicating that the method reliably determines the optimal retrieval amount.
Contriever BGE GTE
fixed-1k 2351.96 2289.33 2265.12
fixed-5k 2236.31 2186.08 2154.81
fixed-10k 2095.38 2056.89 2017.09
fixed-25k 1680.93 1660.64 1599.99
fixed-50k (oracle) 992.47 985.39 898.34
Adaptive-$k$ 1133.28 1248.99 1001.00
Self-Route 2236.31 2186.09 2154.81
Table 4: Diff-$k$ values in the HoloBench task with an information amount of 50k for the three embedding models.
Contriever BGE GTE
fixed-1k (oracle) 162.52 88.86 631.90
fixed-5k 154.68 97.89 598.70
fixed-10k 156.29 116.35 558.69
fixed-25k 191.14 185.70 452.32
fixed-50k 297.41 320.50 307.12
Adaptive-$k$ 161.14 89.40 632.28
Self-Route 154.68 97.89 598.70
Table 5: Diff-$k$ values in HotpotQA for the three embedding models. Given the low retrieval performance with the GTE embeddings in the factoid-type QA tasks in Table 3, the results of GTE embeddings in this table are not informative.
6 Conclusion
We presented a simple yet effective and efficient plug-and-play method, adaptive-$k$, that dynamically selects the number of context chunks to retrieve in a single step, based on the similarity distribution between the query and context chunks. Unlike existing adaptive retrieval methods that requires iterative inference steps, our method only requires a single matrix calculation to estimate the retrieval threshold, achieving a fast and flexible retrieval module. This method is particularly effective for aggregation-type QA tasks, where the optimal number of context chunks varies across examples and cannot be predetermined by a fixed-token retrieval strategy. Results on HoloBench demonstrate that Adaptive-$k$ flexibly adjusts retrieval size to align with the amount of relevant information in the context. In factoid QA tasks, where relevant information is sparse, our method aggressively prunes the context while still outperforming zero-shot QA in answer quality. Compared to Self-Route, our method consistently achieves superior performance in aggregation-type QA tasks, while drastically reducing the input size and maintaining higher context recall.
Our adaptive-$k$ retrieval is a plug-and-play, single-pass alternative to fixed-size retrieval, yet several directions remain. First, because the method is orthogonal to most RAG pipelines, pairing it with techniques such as query-expansion, iterative reranking, or generative feedback loops could further improve accuracy and latency. Second, embedding models excel on different query and corpus traits; a runtime system that selects or ensembles embeddings per query may unlock extra gains in recall and robustness.
Limitations
While our proposed method shows promising results in adaptive retrieval for question answering tasks, it has several limitations that warrant discussion.
[…]
Second, our method is designed for natural language inputs and assumes meaningful semantic similarity between queries and context chunks. It does not generalize well to non-natural-language tasks, such as those involving structured key-value formats (e.g., JSON), where semantic embeddings may not capture relevance effectively.
Third, the approach is sensitive to surface-level variations in text. For example, typographical errors in the query or context can negatively affect embedding quality and distort similarity scores, leading to suboptimal retrieval decisions. If the queries are expected to be noisy with non-standard spellings or grammar, adding a query standardization module Chan et al. (2024) on top of our adaptive-$k$ method would be helpful.
Lastly, the method may be vulnerable to adversarial or malicious inputs Wallace et al. (2019). A specially crafted context chunk could receive an artificially high or low similarity score, thereby introducing a large gap in the similarity distribution and misleading the algorithm into selecting an incorrect retrieval threshold Su et al. (2024). Mitigating such risks would require additional robustness checks or adversarial training techniques, which are beyond the scope of this work.
Ethical considerations
One of the key advantages of our proposed adaptive retrieval method is its potential to reduce the environmental impact of LLM inference. By discarding irrelevant context chunks and only retrieving a minimal yet sufficient subset of documents, our approach significantly reduces the number of input tokens processed. In our experiments, our proposed method discarded nearly 99% of the input tokens in factoid QA tasks, and substantially reduced input size in aggregation QA tasks while maintaining high context recall.
This reduction translates into lower computational overhead, leading to more energy-efficient inference. As a result, our method contributes to decreasing the carbon footprint associated with deploying LLMs at scale. With the growing trend of longer context windows, flexibly filtering out irrelevant context is necessary to ensure energy-efficient inference.
While efficiency is a central goal, we emphasize that any optimization must not compromise fairness or content coverage. Our method is designed to be model-agnostic and does not introduce or amplify biases beyond those present in the similarity scoring mechanism, e.g., cosine similarity over embedding spaces. However, care should be taken when applying this method in high-stakes domains, e.g., medical or legal QA, where discarding seemingly low-similarity context could result in the omission of critical information. Further research is needed to quantify such risks and guide responsible deployment.
While we used AI assitants such as ChatGPT and Copilot to assist in coding and revising this paper, we carefully reviewed and edited all content to ensure it meets our standards and aligns with our research goals.
Acknowledgments
We thank Hayate Iso and Pouya Pezeshkpour for their constructive feedback on this study.
Appendix A Appendix
A.1 Prompt templates
A.1.1 Prompt template for the factoid QA tasks
A.1.2 Prompt template for the HoloBench tasks
A.1.3 Prompt template for LLM-as-a-Judge
A.2 Detailed experimental setup
We set temperature and top-p parameters to $0.0$ and $1.0$, respectively, for all our experiments. For Gemini-2.5-Flash, we set its thinking budget to 0. Table 6 lists the models used in our experiments.
Model Size Context Model name / snapshot License
GPT-4o 128k gpt-4o-2024-08-06 OpenAI Service Terms [^3] https://openai.com/policies/services-agreement/ [Accessed: May 12, 2025]
GPT-4o-mini 128k gpt-4o-mini-2024-07-18 OpenAI Service Terms
Gemini-2.5-Flash 1M gemini-2.5-flash-preview-04-17 Gemini API Additional Terms of Service [^4] https://ai.google.dev/gemini-api/terms [Accessed: May 12, 2025]
Llama-4-Maverick 400B 1M meta-llama/Llama-4-Maverick-17B-128E-Instruct Llama 4 Community License Agreement [^5] https://www.llama.com/llama4/license/ [Accessed: May 12, 2025]
Llama-4-Scout 109B 10M meta-llama/Llama-4-Scout-17B-16E-Instruct Llama 4 Community License Agreement
Table 6: A list of the LLMs used in the experiments. An em-dash (—) means that the model size is not publicly disclosed.
A.3 Full results
A.3.1 Factoid QA tasks (BGE embeddings)
Task Method Score Context recall Reduction (%) $n_{\text{in}}$ $n_{\text{out}}$
HotpotQA zeroshot 39 0.00 ± 0.00 100.00 ± 0.00 63.03 ± 6.98 15.56 ± 9.92
fixed-1k 60 69.33 ± 30.77 99.33 ± 0.09 852.30 ± 57.33 20.53 ± 13.61
fixed-5k 66 84.50 ± 26.93 96.46 ± 0.29 3983.81 ± 219.82 20.87 ± 14.45
fixed-10k 66 88.50 ± 23.05 92.89 ± 0.55 7911.32 ± 440.28 21.85 ± 14.35
fixed-25k 67 92.50 ± 20.15 82.19 ± 1.32 19763.59 ± 1109.32 22.36 ± 14.48
fixed-50k 66 95.33 ± 14.23 64.46 ± 2.45 39597.66 ± 2224.65 24.00 ± 15.87
full-context 45 100.00 ± 0.00 0.00 ± 0.00 109666.92 ± 5537.59 16.49 ± 18.35
self-route 61 90.83 ± 22.77 75.25 ± 40.17 28008.57 ± 45573.55 17.31 ± 17.84
adaptive-k 63 70.83 ± 31.01 99.24 ± 0.17 954.33 ± 206.78 20.46 ± 14.43
NQ zeroshot 49 0.00 ± 0.00 100.00 ± 0.00 53.37 ± 2.29 23.36 ± 15.22
fixed-1k 54 26.45 ± 30.50 99.36 ± 0.09 806.77 ± 70.12 28.50 ± 18.53
fixed-5k 59 42.45 ± 36.68 96.66 ± 0.28 3837.48 ± 333.10 31.33 ± 21.80
fixed-10k 58 50.35 ± 35.94 93.27 ± 0.52 7632.00 ± 655.62 32.11 ± 23.94
fixed-25k 62 62.78 ± 32.87 83.10 ± 1.25 19051.28 ± 1574.22 33.91 ± 26.90
fixed-50k 59 68.89 ± 29.96 66.16 ± 2.45 38102.93 ± 2989.91 37.39 ± 29.33
full-context 41 100.00 ± 0.00 0.00 ± 0.00 110607.54 ± 4711.16 23.10 ± 27.28
self-route 55 52.72 ± 36.37 77.33 ± 38.86 25839.53 ± 44249.67 23.60 ± 20.09
adaptive-k 54 27.20 ± 31.58 99.25 ± 0.25 927.69 ± 283.52 28.83 ± 18.99
TriviaQA zeroshot 83 0.00 ± 0.00 100.00 ± 0.00 60.09 ± 7.79 7.85 ± 6.27
fixed-1k 94 31.21 ± 36.58 99.34 ± 0.10 833.75 ± 67.42 11.86 ± 9.07
fixed-5k 93 42.10 ± 39.93 96.53 ± 0.26 3913.01 ± 260.44 11.73 ± 9.41
fixed-10k 92 49.90 ± 40.11 93.03 ± 0.49 7772.18 ± 488.60 12.64 ± 11.19
fixed-25k 93 54.74 ± 40.26 82.59 ± 1.14 19384.51 ± 1164.94 13.64 ± 11.10
fixed-50k 94 61.66 ± 37.83 65.21 ± 2.27 38819.58 ± 2326.33 15.72 ± 11.95
full-context 61 100.00 ± 0.00 0.00 ± 0.00 110733.69 ± 3419.97 11.95 ± 13.89
self-route 90 48.19 ± 39.84 84.95 ± 31.53 17112.00 ± 35900.97 8.69 ± 8.41
adaptive-k 92 31.21 ± 36.58 99.26 ± 0.23 918.86 ± 240.86 11.69 ± 9.15
Average zeroshot 57.00 0.00 0.00 58.83 15.59
fixed-1k 69.33 42.33 99.34 830.94 20.30
fixed-5k 72.67 56.35 96.55 3911.43 21.31
fixed-10k 72.00 62.92 93.07 7771.83 22.20
fixed-25k 74.00 70.00 82.63 19399.79 23.30
fixed-50k 73.00 75.29 65.28 38840.06 25.70
full-context 49.00 100.00 0.00 110336.05 17.18
self-route 68.67 63.91 79.18 23653.37 16.53
adaptive-k 69.67 43.08 99.25 933.63 20.33
Table 7: Full GPT-4o-mini’s results in the factoid QA tasks.
Task Method Score Context recall Reduction (%) $n_{\text{in}}$ $n_{\text{out}}$
HotpotQA zeroshot 50 0.00 ± 0.00 100.00 ± 0.00 63.03 ± 6.98 19.31 ± 16.53
fixed-1k 61 69.33 ± 30.77 99.33 ± 0.09 852.30 ± 57.33 27.72 ± 18.35
fixed-5k 70 84.50 ± 26.93 96.46 ± 0.29 3983.81 ± 219.82 27.25 ± 17.25
fixed-10k 76 88.50 ± 23.05 92.89 ± 0.55 7911.32 ± 440.28 29.16 ± 18.31
fixed-25k 74 92.50 ± 20.15 82.19 ± 1.32 19763.59 ± 1109.32 28.66 ± 20.86
fixed-50k 73 95.33 ± 14.23 64.46 ± 2.45 39597.66 ± 2224.65 27.89 ± 20.22
full-context 48 100.00 ± 0.00 0.00 ± 0.00 109666.92 ± 5537.59 18.91 ± 20.30
self-route 66 84.50 ± 26.93 96.46 ± 0.29 23663.11 ± 42241.38 22.49 ± 20.04
adaptive-k 63 70.83 ± 31.01 99.24 ± 0.17 954.33 ± 206.78 28.24 ± 19.56
NQ zeroshot 57 0.00 ± 0.00 100.00 ± 0.00 53.37 ± 2.29 27.38 ± 24.23
fixed-1k 59 26.45 ± 30.50 99.36 ± 0.09 806.77 ± 70.12 33.62 ± 26.13
fixed-5k 64 42.45 ± 36.68 96.66 ± 0.28 3837.48 ± 333.10 37.37 ± 30.24
fixed-10k 64 50.35 ± 35.94 93.27 ± 0.52 7632.00 ± 655.62 38.41 ± 32.61
fixed-25k 64 62.78 ± 32.87 83.10 ± 1.25 19051.28 ± 1574.22 38.48 ± 33.35
fixed-50k 63 68.89 ± 29.96 66.16 ± 2.45 38102.93 ± 2989.91 39.76 ± 33.99
full-context 41 100.00 ± 0.00 0.00 ± 0.00 110607.54 ± 4711.16 24.74 ± 33.79
self-route 61 42.45 ± 36.68 96.66 ± 0.28 24713.66 ± 43308.30 33.64 ± 34.15
adaptive-k 61 27.20 ± 31.58 99.25 ± 0.25 927.69 ± 283.52 34.85 ± 27.40
TriviaQA zeroshot 91 0.00 ± 0.00 100.00 ± 0.00 60.09 ± 7.79 8.33 ± 7.16
fixed-1k 96 31.21 ± 36.58 99.34 ± 0.10 833.75 ± 67.42 16.10 ± 11.89
fixed-5k 95 42.10 ± 39.93 96.53 ± 0.26 3913.01 ± 260.44 16.01 ± 11.49
fixed-10k 94 49.90 ± 40.11 93.03 ± 0.49 7772.18 ± 488.60 15.55 ± 9.91
fixed-25k 93 54.74 ± 40.26 82.59 ± 1.14 19384.51 ± 1164.94 15.96 ± 9.81
fixed-50k 93 61.66 ± 37.83 65.21 ± 2.27 38819.58 ± 2326.33 16.31 ± 10.62
full-context 62 100.00 ± 0.00 0.00 ± 0.00 110733.69 ± 3419.97 12.94 ± 12.48
self-route 92 42.10 ± 39.93 96.53 ± 0.26 13900.19 ± 31870.44 11.39 ± 10.31
adaptive-k 96 31.21 ± 36.58 99.26 ± 0.23 918.86 ± 240.86 16.10 ± 12.07
Average zeroshot 66.00 0.00 0.00 58.83 18.34
fixed-1k 72.00 42.33 99.34 830.94 25.81
fixed-5k 76.33 56.35 96.55 3911.43 26.88
fixed-10k 78.00 62.92 93.07 7771.83 27.71
fixed-25k 77.00 70.00 82.63 19399.79 27.70
fixed-50k 76.33 75.29 65.28 38840.06 27.99
full-context 50.33 100.00 0.00 110336.05 18.86
self-route 73.00 56.35 96.55 20758.99 22.51
adaptive-k 73.33 43.08 99.25 933.63 26.40
Table 8: Full GPT-4o’s results in the factoid QA tasks.
Task Method Score Context recall Reduction (%) $n_{\text{in}}$ $n_{\text{out}}$
HotpotQA zeroshot 46 0.00 ± 0.00 100.00 ± 0.00 57.25 ± 7.72 3.00 ± 1.62
fixed-1k 54 69.33 ± 30.77 99.33 ± 0.09 883.55 ± 67.88 15.24 ± 23.19
fixed-5k 63 84.50 ± 26.93 96.46 ± 0.29 4170.89 ± 274.70 12.92 ± 19.22
fixed-10k 66 88.50 ± 23.05 92.89 ± 0.55 8295.62 ± 548.91 12.62 ± 17.17
fixed-25k 66 92.50 ± 20.15 82.19 ± 1.32 20727.06 ± 1379.61 15.09 ± 20.48
fixed-50k 72 95.33 ± 14.23 64.46 ± 2.45 41530.89 ± 2780.34 16.97 ± 22.12
full-context 71 100.00 ± 0.00 0.00 ± 0.00 115121.31 ± 5964.79 17.60 ± 19.32
self-route 68 95.33 ± 17.42 69.45 ± 43.53 36820.47 ± 52663.73 13.95 ± 19.00
adaptive-k 55 70.83 ± 31.01 99.24 ± 0.17 990.43 ± 216.72 15.29 ± 24.47
NQ zeroshot 47 0.00 ± 0.00 100.00 ± 0.00 46.30 ± 2.46 4.67 ± 3.92
fixed-1k 44 26.45 ± 30.50 99.36 ± 0.09 826.53 ± 80.39 26.14 ± 33.25
fixed-5k 59 42.45 ± 36.68 96.66 ± 0.28 3959.77 ± 351.93 31.82 ± 32.77
fixed-10k 59 50.35 ± 35.94 93.27 ± 0.52 7891.38 ± 704.11 115.35 ± 816.66
fixed-25k 62 62.78 ± 32.87 83.10 ± 1.25 19730.89 ± 1697.34 34.87 ± 53.13
fixed-50k 61 68.89 ± 29.96 66.16 ± 2.45 39505.22 ± 3266.78 35.23 ± 47.76
full-context 64 100.00 ± 0.00 0.00 ± 0.00 115142.91 ± 5115.01 28.15 ± 22.01
self-route 60 54.71 ± 36.63 74.41 ± 40.87 30547.34 ± 48902.74 27.40 ± 30.96
adaptive-k 47 27.20 ± 31.58 99.25 ± 0.25 951.38 ± 295.47 29.04 ± 32.84
TriviaQA zeroshot 93 0.00 ± 0.00 100.00 ± 0.00 54.36 ± 8.36 2.55 ± 1.48
fixed-1k 87 31.21 ± 36.58 99.34 ± 0.10 859.44 ± 78.27 8.99 ± 16.82
fixed-5k 93 42.10 ± 39.93 96.53 ± 0.26 4073.06 ± 317.10 8.74 ± 15.25
fixed-10k 92 49.90 ± 40.11 93.03 ± 0.49 8097.22 ± 589.74 9.30 ± 14.94
fixed-25k 93 54.74 ± 40.26 82.59 ± 1.14 20213.64 ± 1424.47 8.41 ± 11.68
fixed-50k 92 61.66 ± 37.83 65.21 ± 2.27 40490.49 ± 2852.03 11.95 ± 14.79
full-context 95 100.00 ± 0.00 0.00 ± 0.00 115669.41 ± 4229.98 10.48 ± 8.73
self-route 94 49.19 ± 39.24 84.95 ± 31.53 17948.59 ± 37786.95 7.39 ± 10.84
adaptive-k 86 31.21 ± 36.58 99.26 ± 0.23 947.50 ± 244.93 7.53 ± 13.74
Average zeroshot 62.00 0.00 0.00 52.64 3.41
fixed-1k 61.67 42.33 99.34 856.51 16.79
fixed-5k 71.67 56.35 96.55 4067.91 17.83
fixed-10k 72.33 62.92 93.07 8094.74 45.76
fixed-25k 73.67 70.00 82.63 20223.86 19.46
fixed-50k 75.00 75.29 65.28 40508.87 21.38
full-context 76.67 100.00 0.00 115311.21 18.74
self-route 74.00 66.41 76.27 28438.80 16.25
adaptive-k 62.67 43.08 99.25 963.10 17.29
Table 9: Full Gemini-2.5-Flash’s results in the factoid QA tasks.
Task Method Score Context recall Reduction (%) $n_{\text{in}}$ $n_{\text{out}}$
HotpotQA zeroshot 38 0.00 ± 0.00 100.00 ± 0.00 61.17 ± 7.05 226.87 ± 1635.13
fixed-1k 65 69.33 ± 30.77 99.33 ± 0.09 850.33 ± 59.50 59.78 ± 97.50
fixed-5k 65 84.50 ± 26.93 96.46 ± 0.29 4006.32 ± 227.71 41.23 ± 63.66
fixed-10k 68 88.50 ± 23.05 92.89 ± 0.55 7963.66 ± 456.87 38.49 ± 60.86
fixed-25k 68 92.50 ± 20.15 82.19 ± 1.32 19912.13 ± 1152.36 33.14 ± 50.95
fixed-50k 67 95.33 ± 14.23 64.46 ± 2.45 39898.83 ± 2310.69 36.25 ± 56.47
full-context 67 100.00 ± 0.00 0.00 ± 0.00 110457.05 ± 5390.30 36.77 ± 65.52
self-route 73 89.50 ± 23.53 84.89 ± 31.51 17292.33 ± 36140.60 60.39 ± 78.22
adaptive-k 63 70.83 ± 31.01 99.24 ± 0.17 952.64 ± 206.19 53.92 ± 87.49
NQ zeroshot 58 0.00 ± 0.00 100.00 ± 0.00 51.54 ± 2.41 62.84 ± 53.57
fixed-1k 62 26.45 ± 30.50 99.36 ± 0.09 802.11 ± 71.62 60.35 ± 48.38
fixed-5k 66 42.45 ± 36.68 96.66 ± 0.28 3847.40 ± 341.79 73.75 ± 67.45
fixed-10k 64 50.35 ± 35.94 93.27 ± 0.52 7660.11 ± 675.79 76.13 ± 61.56
fixed-25k 66 62.78 ± 32.87 83.10 ± 1.25 19136.92 ± 1631.03 85.91 ± 81.62
fixed-50k 68 68.89 ± 29.96 66.16 ± 2.45 38284.28 ± 3103.43 288.08 ± 1664.19
full-context 68 100.00 ± 0.00 0.00 ± 0.00 111154.46 ± 4416.57 113.88 ± 202.07
self-route 66 48.21 ± 36.70 85.06 ± 31.57 17228.44 ± 36372.48 142.10 ± 316.76
adaptive-k 61 27.20 ± 31.58 99.25 ± 0.25 923.58 ± 286.51 67.89 ± 59.55
TriviaQA zeroshot 85 0.00 ± 0.00 100.00 ± 0.00 58.29 ± 7.89 20.12 ± 39.32
fixed-1k 98 31.21 ± 36.58 99.34 ± 0.10 830.54 ± 66.35 24.97 ± 54.24
fixed-5k 98 42.10 ± 39.93 96.53 ± 0.26 3928.14 ± 261.91 24.80 ± 63.48
fixed-10k 97 49.90 ± 40.11 93.03 ± 0.49 7808.89 ± 495.48 18.29 ± 42.28
fixed-25k 96 54.74 ± 40.26 82.59 ± 1.14 19486.58 ± 1189.90 19.47 ± 51.48
fixed-50k 95 61.66 ± 37.83 65.21 ± 2.27 39032.69 ± 2399.86 12.60 ± 22.71
full-context 96 100.00 ± 0.00 0.00 ± 0.00 111322.11 ± 3373.09 23.69 ± 52.77
self-route 97 47.69 ± 40.13 89.78 ± 24.76 11742.38 ± 28544.35 34.38 ± 63.98
adaptive-k 98 31.21 ± 36.58 99.26 ± 0.23 915.89 ± 238.98 30.32 ± 73.20
Average zeroshot 60.33 0.00 0.00 57.00 103.28
fixed-1k 75.00 42.33 99.34 827.66 48.37
fixed-5k 76.33 56.35 96.55 3927.29 46.59
fixed-10k 76.33 62.92 93.07 7810.89 44.30
fixed-25k 76.67 70.00 82.63 19511.88 46.17
fixed-50k 76.67 75.29 65.28 39071.93 112.31
full-context 77.00 100.00 0.00 110977.87 58.11
self-route 78.67 61.80 86.57 15421.05 78.96
adaptive-k 74.00 43.08 99.25 930.70 50.71
Table 10: Full Llama4-Scout’s results in the factoid QA tasks.
Task Method Score Context recall Reduction (%) $n_{\text{in}}$ $n_{\text{out}}$
HotpotQA zeroshot 52 0.00 ± 0.00 100.00 ± 0.00 61.17 ± 7.05 99.43 ± 110.61
fixed-1k 71 69.33 ± 30.77 99.33 ± 0.09 850.33 ± 59.50 110.04 ± 141.75
fixed-5k 78 84.50 ± 26.93 96.46 ± 0.29 4006.32 ± 227.71 68.99 ± 88.75
fixed-10k 74 88.50 ± 23.05 92.89 ± 0.55 7963.66 ± 456.87 44.12 ± 63.11
fixed-25k 71 92.50 ± 20.15 82.19 ± 1.32 19912.13 ± 1152.36 43.30 ± 55.48
fixed-50k 72 95.33 ± 14.23 64.46 ± 2.45 39898.83 ± 2310.69 45.51 ± 66.82
full-context 75 100.00 ± 0.00 0.00 ± 0.00 110457.05 ± 5390.30 41.18 ± 54.88
self-route 79 86.00 ± 26.35 92.61 ± 19.00 8383.42 ± 21450.47 90.55 ± 120.15
adaptive-k 71 70.83 ± 31.01 99.24 ± 0.17 952.64 ± 206.19 112.62 ± 146.43
NQ zeroshot 53 0.00 ± 0.00 100.00 ± 0.00 51.54 ± 2.41 54.22 ± 54.55
fixed-1k 63 26.45 ± 30.50 99.36 ± 0.09 802.11 ± 71.62 68.16 ± 58.41
fixed-5k 65 42.45 ± 36.68 96.66 ± 0.28 3847.40 ± 341.79 75.36 ± 69.75
fixed-10k 67 50.35 ± 35.94 93.27 ± 0.52 7660.11 ± 675.79 75.90 ± 79.63
fixed-25k 64 62.78 ± 32.87 83.10 ± 1.25 19136.92 ± 1631.03 69.41 ± 69.85
fixed-50k 64 68.89 ± 29.96 66.16 ± 2.45 38284.28 ± 3103.43 71.23 ± 77.03
full-context 67 100.00 ± 0.00 0.00 ± 0.00 111154.46 ± 4416.57 66.25 ± 66.57
self-route 65 45.95 ± 37.23 90.86 ± 23.07 10536.57 ± 26530.78 69.87 ± 75.65
adaptive-k 62 27.20 ± 31.58 99.25 ± 0.25 923.58 ± 286.51 75.56 ± 73.03
TriviaQA zeroshot 91 0.00 ± 0.00 100.00 ± 0.00 58.29 ± 7.89 25.52 ± 54.17
fixed-1k 96 31.21 ± 36.58 99.34 ± 0.10 830.54 ± 66.35 39.52 ± 57.18
fixed-5k 98 42.10 ± 39.93 96.53 ± 0.26 3928.14 ± 261.91 31.02 ± 48.63
fixed-10k 98 49.90 ± 40.11 93.03 ± 0.49 7808.89 ± 495.48 21.72 ± 32.93
fixed-25k 96 54.74 ± 40.26 82.59 ± 1.14 19486.58 ± 1189.90 22.79 ± 46.88
fixed-50k 96 61.66 ± 37.83 65.21 ± 2.27 39032.69 ± 2399.86 15.46 ± 34.71
full-context 96 100.00 ± 0.00 0.00 ± 0.00 111322.11 ± 3373.09 19.58 ± 34.15
self-route 98 44.60 ± 40.03 93.64 ± 16.55 7440.28 ± 19973.57 37.87 ± 63.00
adaptive-k 96 31.21 ± 36.58 99.26 ± 0.23 915.89 ± 238.98 49.73 ± 70.11
Average zeroshot 65.33 0.00 0.00 57.00 59.72
fixed-1k 76.67 42.33 99.34 827.66 72.57
fixed-5k 80.33 56.35 96.55 3927.29 58.46
fixed-10k 79.67 62.92 93.07 7810.89 47.25
fixed-25k 77.00 70.00 82.63 19511.88 45.17
fixed-50k 77.33 75.29 65.28 39071.93 44.07
full-context 79.33 100.00 0.00 110977.87 42.34
self-route 80.67 58.85 92.37 8786.76 66.10
adaptive-k 76.33 43.08 99.25 930.70 79.30
Table 11: Full Llama4-Maverick’s results in the factoid QA tasks.
A.3.2 HoloBench (BGE embeddings)
Info amount Method Score Context recall Reduction (%) $n_{\text{in}}$ $n_{\text{out}}$
info5k zeroshot 10.00 0.00 ± 0.00 100.00 ± 0.00 58.13 43.34
fixed-1k 28.19 12.05 ± 6.42 99.18 ± 0.43 1000.07 325.14
fixed-5k 38.74 51.92 ± 29.82 95.85 ± 1.86 4194.01 923.67
fixed-10k 43.50 66.68 ± 30.73 91.80 ± 2.92 8011.64 801.81
fixed-25k 39.81 78.48 ± 26.96 79.59 ± 4.56 19574.94 1489.67
fixed-50k 37.67 86.79 ± 20.88 57.76 ± 5.92 39224.80 2342.40
full-context 37.76 100.00 ± 0.00 0.00 ± 0.00 85882.50 3147.27
self-route 31.32 69.46 ± 27.62 74.68 ± 40.18 23044.97 1523.99
adaptive-k 40.86 75.74 ± 30.48 74.07 ± 25.68 24625.02 2220.94
info10k zeroshot 6.22 0.00 ± 0.00 100.00 ± 0.00 58.13 45.28
fixed-1k 22.55 6.53 ± 3.11 99.19 ± 0.40 1003.83 322.87
fixed-5k 34.06 31.77 ± 14.86 95.84 ± 1.90 4228.71 987.09
fixed-10k 34.85 59.10 ± 27.14 91.74 ± 3.47 8235.02 1712.06
fixed-25k 36.44 78.18 ± 26.55 79.55 ± 5.42 19838.92 1716.79
fixed-50k 29.98 87.34 ± 20.51 57.88 ± 6.41 39437.42 2910.81
full-context 26.59 100.00 ± 0.00 0.00 ± 0.00 86139.74 4370.90
self-route 28.56 48.18 ± 27.16 77.76 ± 37.78 19627.93 2228.59
adaptive-k 33.16 68.54 ± 32.55 79.22 ± 21.59 20233.99 2338.73
info25k zeroshot 4.22 0.00 ± 0.00 100.00 ± 0.00 58.13 43.17
fixed-1k 16.67 2.77 ± 1.15 99.21 ± 0.32 999.62 331.77
fixed-5k 25.76 14.06 ± 5.50 95.96 ± 1.56 4215.72 670.87
fixed-10k 28.61 28.80 ± 9.87 91.86 ± 3.09 8269.06 1554.73
fixed-25k 32.63 68.13 ± 22.77 79.60 ± 7.10 20475.06 2249.00
fixed-50k 30.89 86.88 ± 20.14 58.46 ± 8.58 40152.73 3203.13
full-context 29.53 100.00 ± 0.00 0.00 ± 0.00 86751.63 3767.70
self-route 27.01 34.19 ± 35.59 74.68 ± 40.17 22726.84 1090.31
adaptive-k 25.68 66.16 ± 36.90 73.86 ± 23.04 25778.38 2818.23
info50k zeroshot 5.28 0.00 ± 0.00 100.00 ± 0.00 58.13 43.59
fixed-1k 11.73 1.47 ± 0.55 99.22 ± 0.23 1006.91 310.46
fixed-5k 20.88 7.54 ± 2.52 96.02 ± 1.12 4235.36 511.71
fixed-10k 21.53 15.39 ± 4.25 92.01 ± 2.13 8288.68 1556.82
fixed-25k 25.29 39.55 ± 8.89 79.75 ± 5.38 20622.82 2609.31
fixed-50k 30.18 76.90 ± 16.82 58.85 ± 9.99 41264.63 2885.44
full-context 27.98 100.00 ± 0.00 0.00 ± 0.00 87936.41 3051.19
self-route 19.57 26.38 ± 37.06 76.86 ± 38.66 21444.74 945.72
adaptive-k 22.93 67.43 ± 38.13 60.73 ± 25.94 39654.20 2781.81
Average zeroshot 6.43 0.00 0.00 58.13 43.84
fixed-1k 19.78 5.70 99.20 1002.61 322.56
fixed-5k 29.86 26.32 95.92 4218.45 773.33
fixed-10k 32.12 42.49 91.85 8201.10 1406.36
fixed-25k 33.54 66.09 79.62 20127.94 2016.19
fixed-50k 32.18 84.48 58.24 40019.90 2835.45
full-context 30.47 100.00 0.00 86677.57 3584.26
self-route 26.61 44.55 76.00 21711.12 1447.15
adaptive-k 30.66 69.47 71.97 27572.90 2539.93
Table 12: Full GPT-4o-mini’s results in the HoloBench tasks.
Info amount Method Score Context recall Reduction (%) $n_{\text{in}}$ $n_{\text{out}}$
info5k zeroshot 7.22 0.00 ± 0.00 100.00 ± 0.00 58.13 65.21
fixed-1k 26.14 12.05 ± 6.42 99.18 ± 0.43 1000.07 289.39
fixed-5k 42.32 51.92 ± 29.82 95.85 ± 1.86 4194.01 913.32
fixed-10k 49.79 66.68 ± 30.73 91.80 ± 2.92 8011.64 1334.67
fixed-25k 46.27 78.48 ± 26.96 79.59 ± 4.56 19574.94 2087.61
fixed-50k 43.82 86.79 ± 20.88 57.76 ± 5.92 39224.80 3188.69
full-context 48.30 100.00 ± 0.00 0.00 ± 0.00 73652.50 3680.13
self-route 41.86 65.79 ± 27.76 76.70 ± 38.60 17260.98 1139.29
adaptive-k 48.60 75.74 ± 30.48 74.07 ± 25.68 24625.02 1362.71
info10k zeroshot 5.11 0.00 ± 0.00 100.00 ± 0.00 58.13 64.00
fixed-1k 21.83 6.53 ± 3.11 99.19 ± 0.40 1003.83 298.53
fixed-5k 32.45 31.77 ± 14.86 95.84 ± 1.90 4228.71 1001.83
fixed-10k 36.48 59.10 ± 27.14 91.74 ± 3.47 8235.02 2589.58
fixed-25k 39.65 78.18 ± 26.55 79.55 ± 5.42 19838.92 3527.68
fixed-50k 38.55 87.34 ± 20.51 57.88 ± 6.41 39437.42 4061.82
full-context 41.75 100.00 ± 0.00 0.00 ± 0.00 75768.00 4767.44
self-route 32.37 45.04 ± 25.18 79.96 ± 36.00 17208.18 1146.12
adaptive-k 37.06 68.54 ± 32.55 79.22 ± 21.59 20233.99 2252.20
info25k zeroshot 3.54 0.00 ± 0.00 100.00 ± 0.00 58.13 65.69
fixed-1k 16.51 2.77 ± 1.15 99.21 ± 0.32 999.62 282.18
fixed-5k 27.52 14.06 ± 5.50 95.96 ± 1.56 4215.72 1350.80
fixed-10k 29.10 28.80 ± 9.87 91.86 ± 3.09 8269.06 2520.64
fixed-25k 40.25 68.13 ± 22.77 79.60 ± 7.10 20475.06 3406.40
fixed-50k 34.18 86.88 ± 20.14 58.46 ± 8.58 40152.73 4366.80
full-context 42.24 100.00 ± 0.00 0.00 ± 0.00 75787.37 4802.50
self-route 25.71 30.42 ± 32.70 77.87 ± 37.82 18525.29 1759.81
adaptive-k 33.46 66.16 ± 36.90 73.86 ± 23.04 25778.38 4017.11
info50k zeroshot 5.19 0.00 ± 0.00 100.00 ± 0.00 58.13 62.30
fixed-1k 11.59 1.47 ± 0.55 99.22 ± 0.23 1006.91 274.89
fixed-5k 20.62 7.54 ± 2.52 96.02 ± 1.12 4235.36 735.80
fixed-10k 23.15 15.39 ± 4.25 92.01 ± 2.13 8288.68 1595.42
fixed-25k 28.11 39.55 ± 8.89 79.75 ± 5.38 20622.82 4269.89
fixed-50k 34.58 76.90 ± 16.82 58.85 ± 9.99 41264.63 4244.67
full-context 27.40 100.00 ± 0.00 0.00 ± 0.00 87936.41 3051.19
self-route 19.28 21.54 ± 31.97 80.09 ± 36.03 15426.00 917.82
adaptive-k 30.10 67.43 ± 38.13 60.73 ± 25.94 39654.20 4374.22
Average zeroshot 5.26 0.00 0.00 58.13 64.30
fixed-1k 19.02 5.70 99.20 1002.61 286.25
fixed-5k 30.73 26.32 95.92 4218.45 1000.44
fixed-10k 34.63 42.49 91.85 8201.10 2010.08
fixed-25k 38.57 66.09 79.62 20127.94 3322.89
fixed-50k 37.78 84.48 58.24 40019.90 3965.49
full-context 39.92 100.00 0.00 78286.07 4075.32
self-route 29.80 40.70 78.65 17105.11 1240.76
adaptive-k 37.30 69.47 71.97 27572.90 3001.56
Table 13: Full GPT-4o’s results in the HoloBench tasks.
Info amount Method Score Context recall Reduction (%) $n_{\text{in}}$ $n_{\text{out}}$
info5k zeroshot 10.19 0.00 ± 0.00 100.00 ± 0.00 52.64 644.08
fixed-1k 27.78 12.05 ± 6.42 99.18 ± 0.43 1091.11 478.93
fixed-5k 49.11 51.92 ± 29.82 95.85 ± 1.86 4610.40 1470.04
fixed-10k 54.52 66.68 ± 30.73 91.80 ± 2.92 8760.92 2743.83
fixed-25k 55.31 78.48 ± 26.96 79.59 ± 4.56 21300.94 3422.06
fixed-50k 56.37 86.79 ± 20.88 57.76 ± 5.92 42764.79 3995.68
full-context 63.27 100.00 ± 0.00 0.00 ± 0.00 94227.37 4584.58
self-route 47.56 57.27 ± 31.23 88.38 ± 25.87 11559.12 1894.97
adaptive-k 55.68 75.74 ± 30.48 74.07 ± 25.68 26762.53 2776.24
info10k zeroshot 8.33 0.00 ± 0.00 100.00 ± 0.00 52.64 572.52
fixed-1k 21.09 6.53 ± 3.11 99.19 ± 0.40 1099.17 469.12
fixed-5k 34.94 31.77 ± 14.86 95.84 ± 1.90 4683.90 2107.79
fixed-10k 50.51 59.10 ± 27.14 91.74 ± 3.47 9105.06 2855.68
fixed-25k 55.24 78.18 ± 26.55 79.55 ± 5.42 21717.67 4696.43
fixed-50k 54.06 87.34 ± 20.51 57.88 ± 6.41 43101.23 6040.41
full-context 53.65 100.00 ± 0.00 0.00 ± 0.00 94626.73 6381.93
self-route 35.72 36.72 ± 22.19 89.40 ± 24.10 10196.77 2031.94
adaptive-k 56.26 68.54 ± 32.55 79.22 ± 21.59 22081.99 3637.80
info25k zeroshot 6.28 0.00 ± 0.00 100.00 ± 0.00 52.64 657.34
fixed-1k 15.42 2.77 ± 1.15 99.21 ± 0.32 1098.33 455.64
fixed-5k 31.18 14.06 ± 5.50 95.96 ± 1.56 4699.37 1695.57
fixed-10k 37.86 28.80 ± 9.87 91.86 ± 3.09 9230.02 3053.66
fixed-25k 42.87 68.13 ± 22.77 79.60 ± 7.10 22790.66 6461.89
fixed-50k 44.54 86.88 ± 20.14 58.46 ± 8.58 44317.83 7982.89
full-context 42.19 100.00 ± 0.00 0.00 ± 0.00 95716.53 9289.30
self-route 28.12 20.06 ± 22.11 89.55 ± 24.11 11120.20 2086.64
adaptive-k 43.76 66.16 ± 36.90 73.86 ± 23.04 28207.51 5901.29
info50k zeroshot 6.95 0.00 ± 0.00 100.00 ± 0.00 52.64 735.01
fixed-1k 9.77 1.47 ± 0.55 99.22 ± 0.23 1108.12 445.62
fixed-5k 22.66 7.54 ± 2.52 96.02 ± 1.12 4730.58 1836.44
fixed-10k 30.00 15.39 ± 4.25 92.01 ± 2.13 9277.87 2990.94
fixed-25k 33.71 39.55 ± 8.89 79.75 ± 5.38 23092.23 6596.67
fixed-50k 35.68 76.90 ± 16.82 58.85 ± 9.99 46101.70 8373.53
full-context 45.44 100.00 ± 0.00 0.00 ± 0.00 97597.56 8792.94
self-route 22.37 11.75 ± 19.29 91.76 ± 19.93 8923.41 1610.90
adaptive-k 31.72 67.43 ± 38.13 60.73 ± 25.94 43888.98 7643.86
Average zeroshot 7.94 0.00 0.00 52.64 652.24
fixed-1k 18.51 5.70 99.20 1099.18 462.33
fixed-5k 34.47 26.32 95.92 4681.06 1777.46
fixed-10k 43.22 42.49 91.85 9093.47 2911.03
fixed-25k 46.78 66.09 79.62 22225.38 5294.26
fixed-50k 47.66 84.48 58.24 44071.39 6598.13
full-context 51.14 100.00 0.00 95542.05 7262.19
self-route 33.44 31.45 89.78 10449.88 1906.11
adaptive-k 46.85 69.47 71.97 30235.25 4989.80
Table 14: Full Gemini-2.5-Flash’s results in the HoloBench tasks.
Info amount Method Score Context recall Reduction (%) $n_{\text{in}}$ $n_{\text{out}}$
info5k zeroshot 9.49 0.00 ± 0.00 100.00 ± 0.00 56.18 266.46
fixed-1k 29.52 12.05 ± 6.42 99.18 ± 0.43 994.79 450.14
fixed-5k 40.38 51.92 ± 29.82 95.85 ± 1.86 4195.84 650.72
fixed-10k 40.73 66.68 ± 30.73 91.80 ± 2.92 8024.09 687.39
fixed-25k 37.33 78.48 ± 26.96 79.59 ± 4.56 19625.13 711.01
fixed-50k 34.47 86.79 ± 20.88 57.76 ± 5.92 39342.62 707.11
full-context 36.32 100.00 ± 0.00 0.00 ± 0.00 86093.44 789.47
self-route 36.35 70.18 ± 29.18 69.25 ± 43.22 27700.34 785.69
adaptive-k 39.01 75.74 ± 30.48 74.07 ± 25.68 24711.58 1170.32
info10k zeroshot 7.01 0.00 ± 0.00 100.00 ± 0.00 56.18 260.24
fixed-1k 19.25 6.53 ± 3.11 99.19 ± 0.40 999.90 473.22
fixed-5k 33.58 31.77 ± 14.86 95.84 ± 1.90 4233.97 708.81
fixed-10k 31.87 59.10 ± 27.14 91.74 ± 3.47 8254.73 796.22
fixed-25k 29.75 78.18 ± 26.55 79.55 ± 5.42 19898.58 840.93
fixed-50k 28.60 87.34 ± 20.51 57.88 ± 6.41 39561.99 1070.07
full-context 30.50 100.00 ± 0.00 0.00 ± 0.00 86347.34 1143.99
self-route 33.51 51.87 ± 29.27 72.57 ± 41.54 25307.50 1167.84
adaptive-k 34.69 68.54 ± 32.55 79.22 ± 21.59 20348.23 709.23
info25k zeroshot 7.23 0.00 ± 0.00 100.00 ± 0.00 56.18 277.41
fixed-1k 17.62 2.77 ± 1.15 99.21 ± 0.32 997.53 473.57
fixed-5k 29.44 14.06 ± 5.50 95.96 ± 1.56 4226.03 682.94
fixed-10k 28.95 28.80 ± 9.87 91.86 ± 3.09 8298.52 759.61
fixed-25k 31.39 68.13 ± 22.77 79.60 ± 7.10 20562.19 795.73
fixed-50k 28.35 86.88 ± 20.14 58.46 ± 8.58 40309.13 1268.03
full-context 25.94 100.00 ± 0.00 0.00 ± 0.00 86997.00 961.38
self-route 29.08 32.25 ± 34.30 76.81 ± 38.65 21041.81 914.91
adaptive-k 26.90 66.16 ± 36.90 73.86 ± 23.04 25958.96 854.74
info50k zeroshot 6.18 0.00 ± 0.00 100.00 ± 0.00 56.18 427.23
fixed-1k 11.93 1.47 ± 0.55 99.22 ± 0.23 1004.93 416.12
fixed-5k 22.89 7.54 ± 2.52 96.02 ± 1.12 4246.40 639.41
fixed-10k 23.75 15.39 ± 4.25 92.01 ± 2.13 8317.29 735.40
fixed-25k 23.94 39.55 ± 8.89 79.75 ± 5.38 20706.06 956.62
fixed-50k 25.46 76.90 ± 16.82 58.85 ± 9.99 41427.09 854.82
full-context 22.10 100.00 ± 0.00 0.00 ± 0.00 88197.90 1321.16
self-route 22.24 32.58 ± 40.92 70.50 ± 42.76 27508.79 980.10
adaptive-k 25.45 67.43 ± 38.13 60.73 ± 25.94 39897.64 1182.12
Average zeroshot 7.48 0.00 0.00 56.18 307.84
fixed-1k 19.58 5.70 99.20 999.29 453.26
fixed-5k 31.57 26.32 95.92 4225.56 670.47
fixed-10k 31.33 42.49 91.85 8223.66 744.66
fixed-25k 30.60 66.09 79.62 20197.99 826.07
fixed-50k 29.22 84.48 58.24 40160.21 975.01
full-context 28.72 100.00 0.00 86908.92 1054.00
self-route 30.29 46.72 72.28 25389.61 962.14
adaptive-k 31.51 69.47 71.97 27729.10 979.11
Table 15: Full Llama4-Scout’s results in the HoloBench tasks.
Info amount Method Score Context recall Reduction (%) $n_{\text{in}}$ $n_{\text{out}}$
info5k zeroshot 9.65 0.00 ± 0.00 100.00 ± 0.00 56.18 328.77
fixed-1k 28.05 12.05 ± 6.42 99.18 ± 0.43 994.79 591.14
fixed-5k 48.57 51.92 ± 29.82 95.85 ± 1.86 4195.84 799.80
fixed-10k 54.30 66.68 ± 30.73 91.80 ± 2.92 8024.09 834.79
fixed-25k 56.39 78.48 ± 26.96 79.59 ± 4.56 19625.13 874.41
fixed-50k 53.54 86.79 ± 20.88 57.76 ± 5.92 39342.62 919.02
full-context 55.13 100.00 ± 0.00 0.00 ± 0.00 86093.44 1101.66
self-route 48.47 65.03 ± 29.24 79.91 ± 35.98 18195.21 897.80
adaptive-k 51.77 75.74 ± 30.48 74.07 ± 25.68 24711.58 836.68
info10k zeroshot 9.06 0.00 ± 0.00 100.00 ± 0.00 56.18 322.59
fixed-1k 23.16 6.53 ± 3.11 99.19 ± 0.40 999.90 609.29
fixed-5k 39.02 31.77 ± 14.86 95.84 ± 1.90 4233.97 892.93
fixed-10k 42.91 59.10 ± 27.14 91.74 ± 3.47 8254.73 927.12
fixed-25k 44.54 78.18 ± 26.55 79.55 ± 5.42 19898.58 883.56
fixed-50k 50.37 87.34 ± 20.51 57.88 ± 6.41 39561.99 1020.78
full-context 48.02 100.00 ± 0.00 0.00 ± 0.00 86347.34 1357.50
self-route 37.68 44.92 ± 27.75 79.95 ± 36.00 18010.20 1053.86
adaptive-k 44.91 68.54 ± 32.55 79.22 ± 21.59 20348.23 934.27
info25k zeroshot 7.26 0.00 ± 0.00 100.00 ± 0.00 56.18 322.77
fixed-1k 19.15 2.77 ± 1.15 99.21 ± 0.32 997.53 629.03
fixed-5k 30.36 14.06 ± 5.50 95.96 ± 1.56 4226.03 896.96
fixed-10k 30.28 28.80 ± 9.87 91.86 ± 3.09 8298.52 867.99
fixed-25k 40.47 68.13 ± 22.77 79.60 ± 7.10 20562.19 1053.26
fixed-50k 44.69 86.88 ± 20.14 58.46 ± 8.58 40309.13 1291.90
full-context 43.38 100.00 ± 0.00 0.00 ± 0.00 86997.00 1455.21
self-route 28.57 29.97 ± 33.04 79.00 ± 36.97 19123.37 1038.17
adaptive-k 39.40 66.16 ± 36.90 73.86 ± 23.04 25958.96 1122.81
info50k zeroshot 8.51 0.00 ± 0.00 100.00 ± 0.00 56.18 342.58
fixed-1k 11.24 1.47 ± 0.55 99.22 ± 0.23 1004.93 574.83
fixed-5k 22.39 7.54 ± 2.52 96.02 ± 1.12 4246.40 783.99
fixed-10k 24.53 15.39 ± 4.25 92.01 ± 2.13 8317.29 825.16
fixed-25k 27.94 39.55 ± 8.89 79.75 ± 5.38 20706.06 1127.69
fixed-50k 34.89 76.90 ± 16.82 58.85 ± 9.99 41427.09 1317.22
full-context 36.54 100.00 ± 0.00 0.00 ± 0.00 88197.90 1706.96
self-route 21.90 24.17 ± 35.52 78.99 ± 36.95 19460.93 1183.29
adaptive-k 34.63 67.43 ± 38.13 60.73 ± 25.94 39897.64 1687.78
Average zeroshot 8.62 0.00 0.00 56.18 329.18
fixed-1k 20.40 5.70 99.20 999.29 601.07
fixed-5k 35.09 26.32 95.92 4225.56 843.42
fixed-10k 38.00 42.49 91.85 8223.66 863.76
fixed-25k 42.33 66.09 79.62 20197.99 984.73
fixed-50k 45.87 84.48 58.24 40160.21 1137.23
full-context 45.77 100.00 0.00 86908.92 1405.33
self-route 34.16 41.02 79.46 18697.43 1043.28
adaptive-k 42.68 69.47 71.97 27729.10 1145.38
Table 16: Full Llama4-Maverick’s results in the HoloBench tasks.
A.3.3 Factoid QA tasks (GTE embeddings)
Task Method Score Context recall Reduction (%) $n_{\text{in}}$ $n_{\text{out}}$
HotpotQA zeroshot 50 0.00 ± 0.00 100.00 ± 0.00 63.03 ± 6.98 19.31 ± 16.53
fixed-1k 50 6.00 ± 16.33 99.16 ± 0.20 881.29 ± 76.42 31.19 ± 16.09
fixed-5k 51 9.33 ± 19.44 95.84 ± 0.67 4106.52 ± 327.20 30.61 ± 16.31
fixed-10k 54 13.67 ± 23.73 91.81 ± 1.19 8145.75 ± 576.29 29.29 ± 15.20
fixed-25k 58 27.50 ± 32.17 80.21 ± 2.40 20337.90 ± 1369.28 29.06 ± 17.08
fixed-50k 60 43.50 ± 33.16 61.79 ± 3.81 40520.53 ± 2661.25 28.96 ± 18.42
full-context 48 100.00 ± 0.00 0.00 ± 0.00 109666.92 ± 5537.59 18.91 ± 20.30
self-route 46 25.83 ± 39.17 78.64 ± 37.04 76201.89 ± 52199.49 22.69 ± 20.51
adaptive-k 49 5.50 ± 15.72 99.20 ± 0.36 877.06 ± 409.16 32.87 ± 23.82
NQ zeroshot 57 0.00 ± 0.00 100.00 ± 0.00 53.37 ± 2.29 27.38 ± 24.23
fixed-1k 53 2.65 ± 9.56 99.29 ± 0.19 813.57 ± 84.46 31.61 ± 25.74
fixed-5k 53 9.07 ± 18.72 96.44 ± 0.49 3871.35 ± 353.77 33.67 ± 28.72
fixed-10k 58 14.22 ± 22.67 92.96 ± 0.77 7702.16 ± 668.50 34.24 ± 26.91
fixed-25k 62 31.38 ± 31.59 82.54 ± 1.61 19183.81 ± 1661.81 35.55 ± 29.97
fixed-50k 64 45.18 ± 33.28 65.32 ± 2.88 38288.26 ± 3141.73 36.84 ± 34.84
full-context 41 100.00 ± 0.00 0.00 ± 0.00 110607.54 ± 4711.16 24.74 ± 33.79
self-route 49 20.67 ± 31.77 78.12 ± 38.03 54951.43 ± 55725.18 27.13 ± 25.83
adaptive-k 51 2.85 ± 9.72 99.22 ± 0.24 907.79 ± 263.30 32.75 ± 25.98
TriviaQA zeroshot 91 0.00 ± 0.00 100.00 ± 0.00 60.09 ± 7.79 8.33 ± 7.16
fixed-1k 96 3.00 ± 13.48 99.15 ± 0.49 857.89 ± 96.49 16.98 ± 10.66
fixed-5k 94 4.79 ± 15.73 95.94 ± 0.83 4040.80 ± 358.14 17.35 ± 10.37
fixed-10k 93 6.43 ± 17.79 92.04 ± 1.25 8038.26 ± 670.72 17.64 ± 10.47
fixed-25k 96 16.65 ± 30.03 80.93 ± 1.99 19985.92 ± 1507.53 16.64 ± 10.50
fixed-50k 95 39.57 ± 41.22 63.01 ± 3.13 39806.06 ± 2867.46 17.17 ± 10.11
full-context 62 100.00 ± 0.00 0.00 ± 0.00 110733.69 ± 3419.97 12.94 ± 12.48
self-route 88 10.20 ± 24.97 87.30 ± 27.60 32620.05 ± 48431.50 13.76 ± 10.53
adaptive-k 93 3.00 ± 13.48 99.16 ± 0.56 929.99 ± 604.34 17.33 ± 10.55
Average zeroshot 66.00 0.00 0.00 58.83 18.34
fixed-1k 66.33 3.88 99.20 850.92 26.59
fixed-5k 66.00 7.73 96.07 4006.22 27.21
fixed-10k 68.33 11.44 92.27 7962.06 27.06
fixed-25k 72.00 25.18 81.23 19835.88 27.08
fixed-50k 73.00 42.75 63.38 39538.28 27.66
full-context 50.33 100.00 0.00 110336.05 18.86
self-route 61.00 18.90 81.35 54591.12 21.19
adaptive-k 64.33 3.78 99.19 904.95 27.65
Table 17: Full GPT-4o’s results in the factoid QA tasks with the embeddings by gte-Qwen2-1.5B-instruct.
A.3.4 HoloBench (GTE embeddings)
Info amount Method Score Context recall Reduction (%) $n_{\text{in}}$ $n_{\text{out}}$
info5k zeroshot 7.22 0.00 ± 0.00 100.00 ± 0.00 58.13 65.21
fixed-1k 26.14 12.05 ± 6.42 99.18 ± 0.43 1000.07 289.39
fixed-5k 42.32 51.92 ± 29.82 95.85 ± 1.86 4194.01 913.32
fixed-10k 49.79 66.68 ± 30.73 91.80 ± 2.92 8011.64 1334.67
fixed-25k 46.27 78.48 ± 26.96 79.59 ± 4.56 19574.94 2087.61
fixed-50k 43.82 86.79 ± 20.88 57.76 ± 5.92 39224.80 3188.69
full-context 48.30 100.00 ± 0.00 0.00 ± 0.00 73652.50 3680.13
self-route 40.69 65.18 ± 28.08 76.14 ± 38.33 12113.12 1288.40
adaptive-k 45.23 82.20 ± 25.15 64.79 ± 32.49 29079.80 1879.49
info10k zeroshot 5.11 0.00 ± 0.00 100.00 ± 0.00 58.13 64.00
fixed-1k 21.83 6.53 ± 3.11 99.19 ± 0.40 1003.83 298.53
fixed-5k 32.45 31.77 ± 14.86 95.84 ± 1.90 4228.71 1001.83
fixed-10k 36.48 59.10 ± 27.14 91.74 ± 3.47 8235.02 2589.58
fixed-25k 39.65 78.18 ± 26.55 79.55 ± 5.42 19838.92 3527.68
fixed-50k 38.55 87.34 ± 20.51 57.88 ± 6.41 39437.42 4061.82
full-context 41.75 100.00 ± 0.00 0.00 ± 0.00 75768.00 4767.44
self-route 32.28 45.87 ± 24.77 79.47 ± 35.80 18646.59 1307.47
adaptive-k 39.66 78.99 ± 28.37 65.70 ± 30.96 28475.07 2238.23
info25k zeroshot 3.54 0.00 ± 0.00 100.00 ± 0.00 58.13 65.69
fixed-1k 16.51 2.77 ± 1.15 99.21 ± 0.32 999.62 282.18
fixed-5k 27.52 14.06 ± 5.50 95.96 ± 1.56 4215.72 1350.80
fixed-10k 29.10 28.80 ± 9.87 91.86 ± 3.09 8269.06 2520.64
fixed-25k 40.25 68.13 ± 22.77 79.60 ± 7.10 20475.06 3406.40
fixed-50k 34.18 86.88 ± 20.14 58.46 ± 8.58 40152.73 4366.80
full-context 42.24 100.00 ± 0.00 0.00 ± 0.00 75787.37 4802.50
self-route 23.60 31.02 ± 32.35 77.63 ± 37.70 18905.02 1334.77
adaptive-k 36.33 76.47 ± 31.62 58.43 ± 29.94 36711.21 3509.82
info50k zeroshot 5.19 0.00 ± 0.00 100.00 ± 0.00 58.13 62.30
fixed-1k 11.59 1.47 ± 0.55 99.22 ± 0.23 1006.91 274.89
fixed-5k 20.62 7.54 ± 2.52 96.02 ± 1.12 4235.36 735.80
fixed-10k 23.15 15.39 ± 4.25 92.01 ± 2.13 8288.68 1595.42
fixed-25k 28.11 39.55 ± 8.89 79.75 ± 5.38 20622.82 4269.89
fixed-50k 34.58 76.90 ± 16.82 58.85 ± 9.99 41264.63 4244.67
full-context 27.40 100.00 ± 0.00 0.00 ± 0.00 87936.41 3051.19
self-route 22.08 21.90 ± 31.81 79.88 ± 35.94 17563.86 1843.27
adaptive-k 30.80 72.54 ± 36.80 49.06 ± 29.83 46590.91 3806.21
Average zeroshot 5.26 0.00 0.00 58.13 64.30
fixed-1k 19.02 5.70 99.20 1002.61 286.25
fixed-5k 30.73 26.32 95.92 4218.45 1000.44
fixed-10k 34.63 42.49 91.85 8201.10 2010.08
fixed-25k 38.57 66.09 79.62 20127.94 3322.89
fixed-50k 37.78 84.48 58.24 40019.90 3965.49
full-context 39.92 100.00 0.00 78286.07 4075.32
self-route 29.66 40.99 78.28 16807.15 1443.48
adaptive-k 38.00 77.55 59.49 35214.25 2858.44
Table 18: Full GPT-4o’s results in the HoloBench tasks with the embeddings by gte-Qwen2-1.5B-Instruct.